# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v2

@@ V-D Mitigating Miscompilations
Table VI shows the results of mitigating the Clang miscompilation bugs introduced in Table III with Galápagos. First, we observe that Galápagos successfully generates multiple equivalent variants for each function that triggers each miscompilation bug. Second, Galápagos is able to assemble N-Version binaries for these functions, which caused the execution to crash, instead of producing a wrong output. We consider Galápagos effective in mitigating the three proposed bugs.
@@
TABLE VI: Galápagos’ miscompilation mitigation results.
@@
Figure 11: Control flow graphs for M1. Left: single-version implementation. Right: N-Version implementation.
Figure 12: Disassembled machine code for the hardened M2 function.
@@
{mdframed}
[style=mpdframe,frametitle=Answer to RQ3] Galápagos is able to generate N-Version programs that mitigate the considered real-world miscompilation bugs. Our results show that equivalent function variants produced by LLMs can be employed to harden critical sections of a given program by assembling N-Version functions.
V-E Performance Overhead
Table VII shows a summary of the overhead introduced by the N-Version functions when compared to the reference functions. The column "Baseline Runtime" shows the measurement taken for each function as described in subsection IV-F. For both same-language and cross-language configurations, "N" represents the number of versions used in the maximally diverse N-version function; "N-Version runtime" shows the execution time for the N-version function; and finally "Normalized Overhead" shows the average overhead for a N-version function with one single-variant (i.e. a 2-version function).
@@ V-E Performance Overhead
{mdframed}
[style=mpdframe, frametitle=Answer to RQ4] Our experiments show that the N-Version functions generated by Galápagos introduce the expected overhead. After executing the N-Version functions, we observe that the runtime is increased by 2x for a 2-version function, and grows linearly by adding more variants to the assembly. This experiment demonstrates that the harness overhead overhead is negligible.
Function Name Baseline runtime ($\mu$s) Same-language diversification (C) Cross-language diversification (C+Go)
N N-Version runtime ($\mu$s) Normalized Overhead N N-Version runtime ($\mu$s) Normalized Overhead
@@
TABLE VII: Runtime measurement of N-Version functions.
VI Discussion
@@
VI-A Integration into Existing Software Development Practices
The integration of Galápagos into an existing software development project requires: (1) identifying critical sections of code with the highest reliability standards, and; (2) isolating them into pure functions. Developers would need to integrate Galápagos in the build pipeline for those functions. The pipeline would automatically generate variants, validate their correctness, and insert the N-Version functions into the final build. At deployment time, the N-version functions would terminate execution in case of inconsistencies detected by diversification. This is valid in IoT [38]. We are also aware that companies providing security critical software for the military have such build pipelines with other kinds of diversification.
VI-B Threats to Validity
Threats to internal validity: In our experiments, we use LLVM as the intermediate representation, and alive-tv as the equivalence checker. Therefore, the experimental setting inherits any of their limitations. For instance, some of the variants reported as non-equivalent were not proven as such, but rather the tool failed to process them because of unsupported LLVM instructions.
@@ VI-B Threats to Validity
Threats to external validity: We identify two dimensions where the obtained results can be generalized for Galápagos’ design. First, our experiments focus on two language pairs: C-to-C and C-to-Go. We argue that similar results can be achieved for other language pairs, as long as these are IR-compatible, and a corresponding IR equivalence checking tool exists. Second, our experiments only consider a single LLM as a source of diversity. We argue that similar results can be achieved using different LLMs, as these have been shown to have similar capabilities [18]. We acknowledge that the generalizability of the results, and the claims of suitability of the design can be strengthened by performing experiments with different language pairs, IRs, and generative models.
VII Related Work
VII-A Automated Variant Synthesis
@@
Singh et al. [39] propose the automatic creation of language-diverse, formally verified program variants from a high-level specification. They focus on fault-tolerance of distributed systems by executing each variant following the replicated state machine paradigm. However, no prototype implementation, experimental methodology, or results are provided. Pelofske et al. [18] present a thorough analysis of function variant generation with LLMs. This study is focused on four SHA-1 functions, and provides a deep insight into the diversity and correctness of LLM-generated variant functions. Xu et al. [17] describe a variant generation mechanism based on LLVM-IR obfuscation. Their goal is to create explicitly non-equivalent variants, which are resistant to replication of known tampering mechanisms. While these works aim at automating variant creation, and to different extents, correctness verification, the key novelty of our work is the focus on hardening of binaries via N-Version function execution.
VII-B LLMs code correctness
@@
Pan et al. [40] identify the correctness of LLM-generated code as an important challenge to address. In this vein, Kessel and Atkinson [41] present LASSO, a platform for differential execution of large sets of code versions and tests. Compared to this work Galápagos goes one step further and automates the creation of N-Version programs, in addition to generation and testing of code variants. Du et al. [42] present a benchmark for solutions to programming tasks via LLMs. This benchmark is designed to measure how efficient are the generated solutions in terms of algorithmic complexity. The experiments show that LLMs are able to produce code that solves the tasks while having diverse efficiency scores. Zhang et al. [28] introduce a benchmark to quantify the correctness of code produced by an LLM. The experiments evidence that LLMs have the capability to correctly solve programming tasks to a large extent. However, while the data provided by these works gives a detailed overview of the correctness and efficiency of code produced by LLMs, the evaluations are test-based, and not formally verified against a specification.
VII-C Code Generation and Code Translation
@@
Code generation and translation with LLMs are very active research areas [43, 44]. Yang et al. [22] present a C-to-Rust verified translation tool. It works by using static methods to construct a known equivalent translation, and LLMs to create more idiomatic translations. The latter are then checked for equivalence against the former using an off-the-shelf Rust tool. Bathia et al. [45] describe a tool to produce verified code transpilations. This is achieved by leveraging intermediate representations and formal equivalence checkers. Yang et al. [46] present UniTrans, a framework to generate and test LLM code translation. UniTrans uses the original code to further generate test cases for the resulting translation, and leverage them to iteratively repair the translation in case of failure. While the mentioned works address the issue of validating correctness of LLM-generated code equivalence, their scope is not concerned with software reliability.
VII-D Code Transplantation and Recombination:
@@
As part of Galápagos’ Harnessing pass, IR code from newly generated programs is inserted into the project that is being hardened. This process is similar to the code transplantation mechanisms described in the literature. This process has been studied in the context of automated program repair [47], feature transfer between systems [48, 49, 50, 51], and code clone analysis [52]. The Harnessing pass is also similar to recombination approaches, where different versions of systems are created by manipulating sections of readily available codebases, at lower-than-source levels [53, 54]. While closely related, these works are not concerned with the challenges of the two first passes of Galápagos, namely, automatic diversification and correctness checking.
VII-E Addressing Miscompilations
@@
The works of Eide et al. [26] and Le et al. [55] present approaches to discover miscompilation bugs via stochastic exploration of the program space. With the same goal, Tu et al. [56] describe a mechanism with guided LLM prompt production approach, which allows producing bug-triggering program mutations. While these approaches seek to actively discover miscompilation bugs, Galápagos takes a program hardening, as it does not target compilers directly, but protects critical application code against potential miscompilation.
VIII Conclusion
In this paper, we have introduced the original idea of automated N-Version programming with LLMs. The key idea is to exploit the search power and coding capabilities of LLMs, within the scope of N-Version programming. We presented Galápagos, a system that allows for automatic N-Version programming, with formal guarantees. We evaluated it in four dimensions: variant correctness, code diversity, the ability to counteract miscompilation bugs, and introduced overhead.
Our experiments show that: (1) Galápagos is able to produce code variants with LLMs, that can be proven to be equivalent, even when these variants are written in a different programming language. (2) Galápagos is able to produce code variants which are diverse both on disk and at execution-time; (3) the Galápagos variants can be harnessed to harden critical sections of software against the dangerous class of miscompilation bugs, and; (4) the N-Version functions generated by Galápagos introduce a performance overhead that stays within reasonable and expected limits. Finally, our experiments clearly show that the production of N-Version software can be automated by the combined use of off-the-shelf tools and generative language models.
We identify two future angles for this novel research direction. Galápagos relies on N-of-N voting, which means that all variants must produce the same output for the program to continue to execute. However, previous works in the N-Version programming line have proposed alternative voting mechanisms [57]. Galápagos can be extended to support different voting mechanisms, so that voting can be tailored to the context and specific requirements of the application. For instance, a threshold voting mechanism (k-of-N) would allow Galápagos to provide similar hardening enhancements, while also hiding unwanted behavior, and continue with execution.
Effective prompt engineering is also crucial in guiding LLMs to generate the desired code variants. For instance, future prompts could include the function’s purpose, input and output specifications, constraints, and examples of desired code patterns. A promising direction is to enrich the prompt with measurements of previously generated variants, such as global alignment, Hamming distance, or Levenshtein distance [58, 59, 60].
Finally, we envision LLMs being fine-tuned for diversity. A fine-tuning loop would help produce code with higher degrees of variant diversity and correctness, as it has been for program repair [61]. Fine-tuning could involve adapting pre-trained models to specific diversification tasks or language pairs, in order to enhance their ability to generate diverse and correct code variants.

### v3

@@ 5.4. Mitigating Miscompilations
Table 7 shows the results of mitigating the Clang miscompilation bugs introduced in Table 3 with Galápagos. First, we observe that Galápagos successfully generates multiple equivalent variants for each function that triggers each miscompilation bug. Second, Galápagos is able to assemble N-Version binaries for these functions, which caused the execution to crash, instead of producing a wrong output. We consider Galápagos effective in mitigating the three proposed bugs.
@@
Table 7. Galápagos’ miscompilation mitigation results.
@@
Figure 11. Control flow graphs for M1. Left: single-version implementation. Right: N-Version implementation.
Figure 12. Disassembled machine code for the hardened M2 function.
@@
Galápagos is able to generate N-Version programs that mitigate the considered real-world miscompilation bugs. Our results show that equivalent function variants produced by LLMs can be employed to harden critical sections of a given program by assembling N-Version functions.
5.5. Performance Overhead
Table 8 shows a summary of the overhead introduced by the N-Version functions when compared to the reference functions. The column "Baseline Runtime" shows the measurement taken for each function as described in subsection 4.6. For both same-language and cross-language configurations, "N" represents the number of versions used in the maximally diverse N-Version function; "N-Version runtime" shows the execution time for the N-Version function; and finally "Normalized Overhead" shows the average overhead for a N-Version function with one single-variant (i.e. a 2-version function).
@@ 5.5. Performance Overhead
Our experiments show that the N-Version functions generated by Galápagos introduce the expected overhead. After executing the N-Version functions, we observe that the runtime is increased in average by 2x for a 2-version function, and grows linearly by adding more variants to the assembly. This experiment provides evidence that Galápagos’s harness pass’ overhead is negligible. We note that variants may introduce algorithmic differences in the diversified function (e.g. iteration vs recursion) changing the diversified function complexity itself, yet this does not change the overhead due to the N-Version assembly itself.
[…]
@@
Table 8. Runtime measurement of N-Version functions.
6. Discussion
@@
6.1. Integration into Existing Software Development Practices
The integration of Galápagos into an existing software development project requires: (1) identifying critical sections of code with the highest reliability standards, and; (2) isolating them into pure functions. Developers would need to integrate Galápagos in the build pipeline for those functions. The pipeline would automatically generate variants, validate their correctness, and insert the N-Version functions into the final build. At deployment time, the N-Version functions would terminate execution in case of inconsistencies detected by diversification. This is valid in IoT (almohri2022dynamic, ). We are also aware that companies providing security critical software for the military have such build pipelines with other kinds of diversification.
6.2. Threats to Validity
Threats to internal validity: Per our design, we rely on existing intermediate representations (LLVM IR) and equivalence checkers (alive-tv). Therefore, Galápagos will inherit their trustworthiness and limitations. For instance, in our experiments, some of the variants soundly discarded as non-equivalent were not proven as such, but alive-tv failed to process them because of unsupported LLVM instructions.
@@ 6.2. Threats to Validity
Threats to external validity: We identify two dimensions where the obtained results can be generalized for Galápagos’ design. First, our experiments focus on two language pairs: C-to-C and C-to-Go. We argue that similar results can be achieved for other language pairs, as long as these are IR-compatible, and a corresponding IR equivalence checking tool exists. Second, our experiments only consider a single LLM as a source of diversity. We argue that similar results can be achieved using different LLMs, as these have been shown to have similar capabilities (los-alamos, ). We acknowledge that the generalizability of the results, and the claims of suitability of the design can be strengthened by performing experiments with different language pairs, IRs, generative models, and prompt templates.
7. Related Work
7.1. Automated Variant Synthesis
@@
Singh et al. (avatars, ) propose the automatic creation of language-diverse, formally verified program variants from a high-level specification. They focus on fault-tolerance of distributed systems by executing each variant following the replicated state machine paradigm. However, no prototype implementation, experimental methodology, or results are provided. Pelofske et al. (los-alamos, ) present a thorough analysis of function variant generation with LLMs. This study is focused on four SHA-1 functions, and provides a deep insight into the diversity and correctness of LLM-generated variant functions. Xu et al. (n-version-obf, ) describe a variant generation mechanism based on LLVM-IR obfuscation. Their goal is to create explicitly non-equivalent variants, which are resistant to replication of known tampering mechanisms. While these works aim at automating variant creation, and to different extents, correctness verification, the key novelty of our work is the focus on hardening of binaries via N-Version function execution.
7.2. LLMs code correctness
@@
Pan et al. (survey-bugs, ) identify the correctness of LLM-generated code as an important challenge to address. In this vein, Kessel and Atkinson (Kessel_2024, ) present LASSO, a platform for differential execution of large sets of code versions and tests. Compared to this work Galápagos goes one step further and automates the creation of N-Version programs, in addition to generation and testing of code variants. Du et al. (mercury, ) present a benchmark for solutions to programming tasks via LLMs. This benchmark is designed to measure how efficient are the generated solutions in terms of algorithmic complexity. The experiments show that LLMs are able to produce code that solves the tasks while having diverse efficiency scores. Zhang et al. (NEURIPS2023_43e9d647, ) introduce a benchmark to quantify the correctness of code produced by an LLM. The experiments evidence that LLMs have the capability to correctly solve programming tasks to a large extent. However, while the data provided by these works gives a detailed overview of the correctness and efficiency of code produced by LLMs, the evaluations are test-based, and not formally verified against a specification.
7.3. Code Generation and Code Translation
@@
Code generation and translation with LLMs are very active research areas (survey1, ; survey2, ). Yang et al. (vert, ) present a C-to-Rust verified translation tool. It works by using static methods to construct a known equivalent translation, and LLMs to create more idiomatic translations. The latter are then checked for equivalence against the former using an off-the-shelf Rust tool. Bathia et al. (llmlift, ) describe a tool to produce verified code transpilations. This is achieved by leveraging intermediate representations and formal equivalence checkers. Yang et al. (unitrans, ) present UniTrans, a framework to generate and test LLM code translation. UniTrans uses the original code to further generate test cases for the resulting translation, and leverage them to iteratively repair the translation in case of failure. While the mentioned works address the issue of validating correctness of LLM-generated code equivalence, their scope is not concerned with software reliability.
7.4. Code Transplantation and Recombination:
@@
As part of Galápagos’ Harnessing pass, IR code from newly generated programs is inserted into the project that is being hardened. This process is similar to the code transplantation mechanisms described in the literature. This process has been studied in the context of automated program repair (semgraft, ), feature transfer between systems (carbon-copy, ; plastic-surgery, ; babel-pidgin, ), and code clone analysis (code-clones, ). The Harnessing pass is also similar to recombination approaches, where different versions of systems are created by manipulating sections of readily available codebases, at lower-than-source levels (frankenssl, ; object-level-recombination, ). While closely related, these works are not concerned with the challenges of the two first passes of Galápagos, namely, automatic diversification and correctness checking.
7.5. Addressing Miscompilations
@@
The works of Eide et al. (regehr-volatile, ) and Le et al. (miscomp-guided, ) present approaches to discover miscompilation bugs via stochastic exploration of the program space. With the same goal, Tu et al. (miscomp-llm, ) describe a mechanism with guided LLM prompt production approach, which allows producing bug-triggering program mutations. While these approaches seek to actively discover miscompilation bugs, Galápagos takes a program hardening, as it does not target compilers directly, but protects critical application code against potential miscompilation.
8. Future Work
Looking ahead, we see concrete directions opened by this research.
8.1. Independence of LLM-Generated Variants
A fundamental challenge in N-Version programming is ensuring that different implementations fail independently under the same inputs. While our current approach generates variants that are statically and dynamically different, we need to explore how to measure independence of implementations. This includes defining quantitative metrics for independence, establishing experimental protocols to estimate fault correlation under diverse inputs. Then, one could use these metrics to synthesize variants that actively maximize independence. Future work should also investigate whether LLM-generated variants exhibit different failure patterns, not only between themselves, but compared to human-written implementations.
8.2. LLM Prompting and Search Approaches
The current diversification approach uses a single-shot prompting strategy. Future work should explore more sophisticated approaches to variant generation. This includes casting variant synthesis as constrained search that maximizes behavioral diversity under an equivalence oracle, and exploiting verifier counterexamples to repair failing variants. We should also investigate measurement-driven synthesis, where prompts are conditioned on diversity metrics (e.g., edit distance, instruction-set Jaccard) to steer generation away from previously explored regions. Additionally, diversity-aware fine-tuning could adapt LLMs to the diversification task, with training objectives that encourage behavioral distance while preserving equivalence.
8.3. Language Support Extension
While our current implementation focuses on C and Go, extending Galápagos to support additional programming languages presents both opportunities and challenges. The primary technical barrier is the requirement for IR compatibility—languages must compile to a common intermediate representation (e.g., LLVM IR) to enable formal equivalence checking. This constraint limits support to languages with mature LLVM front-ends, such as Rust or Swift. The implementation cost for adding a new language pair involves: (1) ensuring robust compilation to LLVM IR, (2) adapting the validation pipeline to handle language-specific IR patterns, and (3) potentially developing language-specific equivalence checking tools if existing ones are insufficient. However, the modular design of Galápagos’s three-pass architecture means that language support can be added incrementally without major architectural changes. The diversification pass is already language-agnostic, requiring only appropriate prompting strategies. The validation pass depends on IR-level tools that can be extended or replaced as needed. The harnessing pass operates at the IR level and is largely language-independent. Thus, while each new language pair requires careful engineering effort, the overall approach scales well to additional languages.
9. Conclusion
We introduced automated N-Version programming powered by LLMs and presented Galápagos, a practical pipeline that (1) diversifies functions at the source level, (2) validates candidates via compilation, testing, and formal IR-level equivalence checking, and (3) harnesses verified variants into executable N-Version software.
Across 30 real-world C functions, we found that LLMs can synthesize function variants that are provably equivalent, even across languages. We showed that a substantial portion of the induced diversity survives compiler optimizations and can be measured at runtime. We also demonstrated that assembling these variants as N-Version functions mitigates real miscompilation bugs, by converting silent wrong results into explicit program halts. The added runtime cost matches expectations ($\sim$2x for a 2-version function) and scales linearly with N. Overall, our results indicate that combining practical formal methods with modern LLMs can substantially reduce the engineering cost of N-Version programming while giving significant correctness guarantees.